# LLM-assisted Explicit and Implicit Multi-interest Learning Framework for Sequential Recommendation

### v1

@@ 3.4.3. Recommendation Task
For the main recommendation task in the training phase, we maintain the same operation as the traditional SR models. Specifically, the score of each candidate item is obtained by multiplying the user behavior interest representation and the item representation, and then softmax is applied to convert it into probability.
(13) $\hat{y_{i}^{k}}=\text{softmax}(h_{\text{ex}}^{\top}e_{i}^{k}),$
where $e_{i}^{k}$ is item embeddings.
(14) $\mathcal{L}_{R}=\sum_{k=1}^{n}y_{i}^{k}log(\hat{y_{i}^{k}}),$
where $y_{i}^{k}$ is ground truth. Finally, We combine the auxiliary tasks (semantic prediction, modality alignment) with the main task (recommendation) through joint learning to obtain the final loss function.
@@ 4.1.1. Datasets and Evaluation Metrics
Datasets Users Items Interactions Avg.len.
Beauty 22,363 12,101 198,502 8.87
Grocery 14,681 8,713 151,254 10.30
Office 4,905 2,420 53,258 10.85
Table 2. The result of performance comparison between baselines and EIMF on large datasets (Grocery and Beauty).
Dataset Model Pop DNN GRU4Rec SASRec Bert4Rec LLM2Bert4Rec MIND ComiRec-SA REMI EIMF(REMI) Improv.(%)
Grocery $\text{Recall}@20$ 0.0729 0.1252 0.1387 0.1536 0.1544 0.1259 0.1445 0.1122 0.1617 0.1758 +8.71
$\text{Recall}@50$ 0.1305 0.2044 0.2300 0.2508 0.2372 0.2080 0.2162 0.2076 0.2574 0.2704 +5.05
$\text{NDCG}@20$ 0.0448 0.0804 0.0905 0.1041 0.1074 0.0793 0.0844 0.0706 0.0953 0.1025 -4.56
$\text{NDCG}@50$ 0.0624 0.0969 0.1097 0.1139 0.1130 0.0954 0.0967 0.0918 0.1108 0.1181 +3.69
$\text{HR}@20$ 0.1252 0.2076 0.2321 0.2586 0.2614 0.2164 0.2328 0.1851 0.2539 0.2750 +6.34
$\text{HR}@50$ 0.2137 0.3227 0.3649 0.3750 0.3608 0.3294 0.3322 0.3220 0.3832 0.3955 +3.21
@@
To evaluate the framework’s performance, we selected three sub-datasets from the review dataset of the Amazon e-commerce platform [^1] https://jmcauley.ucsd.edu/data/amazon/links.html, namely Office Products, Grocery and Gourmet Food, and All Beauty. The statistics of the three datasets after preprocessing are summarized in Table 1.
In terms of metrics, we chose the Recall @$K$, Normalized Discounted Cumulative Gain (NDCG) @$K$, and Hit Ratio (HR) @$K$, which are commonly used in SR. $K$ value is set to 20 and 50.
4.1.2. Implementation Details and Hyperparameter Settings
The Qwen-Turbo model (Bai et al., 2023) was chosen for the LLM in the experiments and is responsible for inference about user interests. All models were implemented in Pytorch 1.10.0 and Python 3.9 in the conda environment. We follow previous research(Xie et al., 2023), and in our experiments, the batch size $=128$, the dimension $=64$, and the maximum number of training iterations for all models is 1 million. The number of interests for the multi-interest model is set to 4, the number of user interests inferred by LLM to at most $20$ and the optimizer is trained using Adam with the learning rate set to $0.001$. The hyper-parameters $\alpha$ and $\beta$ in the Modal Representation Alignment task were set to 0.4 and 0.1, respectively, and the loss function in the auxiliary task parameter $\gamma$ was set to 0.1. Please note that except for the hyperparameter experiments, we use $p$ = -10 as the reference value setting for the AP algorithm in the rest of the experiments (clustered typical samples account for about 3% of the dataset).
@@ 4.1.3. Baselines
Regarding baseline selection, we mainly select from single-interest and multi-interest aspects, including single-interest models (Covington et al., 2016; Hidasi, 2015; Kang and McAuley, 2018; Sun et al., 2019), LLM-based RS model (Harte et al., 2023), and multi-interest models (Li et al., 2019; Cen et al., 2020; Xie et al., 2023).
  • Pop. An algorithm that makes recommendations based on product popularity.
  • DNN (Covington et al., 2016). The twin-tower DNN model developed by the YouTube team pools user behaviors and then uses MLP to obtain user interest representations.
@@
  • Bert4Rec (Sun et al., 2019). An SR model based on the Bert architecture, which uses pre-trained language model technology to process user behavior sequences and captures contextual information in the sequence through a bidirectional encoder.
  • LLM2Bert4Rec (Harte et al., 2023). A framework to enhance existing SR models by leveraging semantically rich item representations provided by LLM.
@@ 4.2.1. Performance comparison experiment with baselines
Table 3. The result of performance comparison between baselines and EIMF on small dataset (Office).
Model $\text{Recall}@20$ $\text{NDCG}@20$ $\text{HR}@20$ $\text{Recall}@50$ $\text{NDCG}@50$ $\text{HR}@50$
GRU4Rec 0.0800 0.0537 0.1384 0.1770 0.0809 0.2811
SASRec 0.1152 0.0643 0.1792 0.1963 0.0805 0.2871
Bert4Rec 0.1037 0.0635 0.1812 0.1920 0.0834 0.3116
MIND 0.0915 0.0513 0.1466 0.1593 0.0706 0.2505
ComiRec-SA 0.0788 0.0456 0.1202 0.1589 0.0653 0.2321
REMI 0.1072 0.0637 0.1751 0.2018 0.0825 0.2973
LLM2Bert4Rec 0.1085 0.0660 0.1812 0.2117 0.0939 0.3360
EIMF(REMI) 0.1222 0.0649 0.1772 0.2296 0.0930 0.3503
Improv.(%) +6.08 -1.66 -2.20 +8.45 -0.95 +4.25
Table 4. Performance experimental results of different backbones on Beauty and Office datasets.
Dataset Beauty Office
Model $\text{Recall}@50$ $\text{NDCG}@50$ $\text{HR}@50$ $\text{Recall}@50$ $\text{NDCG}@50$ $\text{HR}@50$
Bert4Rec 0.2050 0.0852 0.2986 0.1920 0.0834 0.3116
EIMF(Bert4Rec) 0.2155 0.0854 0.3071 0.2101 0.0909 0.3380
Improv.(%) +5.12 +0.23 +2.84 +9.42 +8.99 +8.47
SASRec 0.2220 0.0879 0.3129 0.1964 0.0806 0.2872
EIMF(SASRec) 0.2499 0.0999 0.3469 0.2088 0.0878 0.3198
Improv.(%) +12.56 +13.65 +10.86 +6.31 +8.93 +11.35
MIND 0.2384 0.0900 0.3272 0.1594 0.0707 0.2505
EIMF(MIND) 0.2518 0.0968 0.3424 0.2023 0.0793 0.2953
Improv.(%) +5.62 +7.55 +4.64 +26.91 +12.16 +17.88
@@
We added the EIMF framework to REMI to conduct performance comparison experiments with other baselines. The results are shown in Table 2. Here we marked the best performance value among all models in bold and the best performance value in the baseline with an underline. Based on the results in Table 2 and Table 3, we can make the following observations:
@@
  • In the large-scale data set in Table 2, the performance of the multi-interest model is significantly better than that of the single-interest model, and as the amount of data increases, this advantage becomes more and more significant. Our analysis shows that multi-interest learning is more suitable for data-rich environments because it is able to capture more complex and diverse underlying interests among users or items, thereby more accurately understanding subtle differences between individuals.
@@
  • In the small-scale dataset shown in Table 3, the traditional single-interest SR model outperforms the multi-interest model. Our analysis shows that when the amount of data is limited, it is difficult to accurately capture the diverse interests of users by simply relying on user behavior information. Therefore, a model that focuses on a single major interest can utilize the limited information in a more concentrated manner, thus providing more accurate recommendation results.
@@
  • In the large datasets, EIMF(REMI) achieved the best performance; while in the small dataset, LLM2Bert4Rec and EIMF(REMI) performed best. First, this shows that the rich semantic knowledge in the LLM can effectively enhance the behavioral modeling ability of user representation, showing the great potential of LLM in recommendation tasks. Second, although some metrics of EIMF(REMI) are slightly lower than LLM2Bert4Rec on small datasets, its better performance on large datasets proves that EIMF(REMI) not only has good stability but also can adapt to more diverse data environments.

### v3

@@ 3.5.3. Recommendation Task
For the main recommendation task in the training phase, we maintain the same operation as the traditional SR models. Specifically, the score of each candidate item is obtained by multiplying the user behavior interest representation and the item representation, and then a softmax is applied to convert it into a probability.
(13) $\hat{y_{k}}=\text{softmax}({{\mathbf{H}_{\text{im}}^{{}^{\prime}}}}^{\top}\mathbf{e}_{k}),$
where $\mathbf{e}_{k}\in\mathbf{E}$ is item embeddings, $\hat{y_{k}}=\mathbb{R}^{N}$ represents the probability of the item appearing as the next click of the user.
(14) $\mathcal{L}_{R}=\sum_{k=1}^{N}y_{k}log(\hat{y_{k}}),$
where $y_{k}$ is the ground truth of candidate item $i_{k}$. Finally, We combine the auxiliary tasks (semantic prediction, modality alignment) with the main task (recommendation) through joint learning to obtain the final loss function.
@@
4. Training Phase Time Complexity Analysis
In this section, we analyze the time complexity of the training phase of EIMF, which encompasses three main steps: precomputing semantic representations with ESIM, processing behavioral data with IBIM, and jointly optimizing three tasks—recommendation, semantic prediction, and modality alignment. We detail the time complexity analysis for each component below. Notations are as defined in Table 1, with additions: $I$ (AP clustering iterations), $d_{l}$ (LLM hidden dimension), $P$ (model parameters), and $E$ (training epochs).
Table 2. Time Complexity Analysis of the EIMF Training Phase
Phase/Component Sub-step Time Complexity Dominant Term
ESIM - Precomputation Text Encoding $O(|\mathcal{U}|Ld_{t})$ $O(|\mathcal{U}|Ld_{t})$
AP Clustering $O(I|\mathcal{U}|^{2})$ $O(I|\mathcal{U}|^{2})$
Constructing Typical Prompts $O(GL)$ $O(GL)$
Inferring Semantic Interests $O(GLd_{l}P)$ $O(GLd_{l}P)$
Semantic Representation Refinement $O(G(M_{\text{ex}}^{2}d_{t}+M_{\text{ex}}d_{t}))$ $O(GM_{\text{ex}}^{2}d_{t})$
Total ESIM $O(|\mathcal{U}|Ld_{t}+I|\mathcal{U}|^{2}+GLd_{l}P+GM_{\text{ex}}^{2}d_{t}+GM_{\text{ex}}d_{t})$ $O(I|\mathcal{U}|^{2})$
IBIM Multi-interest SR Model $O(|\mathcal{U}|L^{2}d)$ $O(|\mathcal{U}|L^{2}d)$
Target-Aware Attention $O(|\mathcal{U}|M_{\text{im}}d)$ $O(|\mathcal{U}|M_{\text{im}}d)$
Recommendation Score $O(|\mathcal{U}|Nd)$ $O(|\mathcal{U}|Nd)$
Total IBIM $O(|\mathcal{U}|(L^{2}d+M_{\text{im}}d+Nd))$ $O(|\mathcal{U}|Nd)$
Semantic Prediction Task Attention Mechanisms $O(|\mathcal{U}|(M_{\text{ex}}^{2}d_{t}+M_{\text{ex}}d_{t}))$ $O(|\mathcal{U}|M_{\text{ex}}^{2}d_{t})$
Scoring and Loss $O(|\mathcal{U}|Nd_{t})$ $O(|\mathcal{U}|Nd_{t})$
Total $O(|\mathcal{U}|(M_{\text{ex}}^{2}d_{t}+M_{\text{ex}}d_{t}+Nd_{t}))$ $O(|\mathcal{U}|Nd_{t})$
Modality Alignment Task Loss Computation $O(|\mathcal{U}|(M_{\text{im}}d+M_{\text{ex}}d_{t}+Nd))$ $O(|\mathcal{U}|Nd)$
Total $O(|\mathcal{U}|(M_{\text{im}}d+M_{\text{ex}}d_{t}+Nd))$ $O(|\mathcal{U}|Nd)$
Recommendation Task Scoring $O(|\mathcal{U}|Nd)$ $O(|\mathcal{U}|Nd)$
Total $O(|\mathcal{U}|Nd)$ $O(|\mathcal{U}|Nd)$
Joint Learning Backpropagation $O(EP|\mathcal{U}|)$ $O(EP|\mathcal{U}|)$
Total $O(EP|\mathcal{U}|)$ $O(EP|\mathcal{U}|)$
Total Training Complexity $O(I|\mathcal{U}|^{2}+E(|\mathcal{U}|Nd+|\mathcal{U}|Nd_{t}+P|\mathcal{U}|))$ $O(I|\mathcal{U}|^{2}+E|\mathcal{U}|Nd)$
The dominant terms are $O(I|\mathcal{U}|^{2})$ from AP clustering, which is a bottleneck for large $|\mathcal{U}|$ but occurs once, and $O(E|\mathcal{U}|Nd)$ from the recommendation task, which scales with training epochs. In the EIMF framework, the ESIM operates exclusively during the training phase through a one-time offline procedure. By clustering users’ textualized interaction sequences and selecting $G\ll|\mathcal{U}|$ representative samples, the original LLM inference complexity of $O(|\mathcal{U}|Ld_{l}P)$ is reduced to $O(GLd_{l}P)$. In practice, representative samples account for less than 3% of the total data, significantly reducing both the computational cost and inference time associated with LLM usage. This design shifts the large-scale LLM inference task to the offline phase, avoiding high latency during online recommendation and laying a solid foundation for low-latency, high-throughput serving in the inference phase.
For the IBIM, EIMF adopts a mature multi-interest SR model to capture users’ behavioral patterns. This module maintains linear scalability with respect to the number of users. After incorporating explicit semantic features, user interest representations are enhanced with minimal additional inference complexity, maintaining high efficiency during serving. To further improve the representational capacity, EIMF introduces two auxiliary tasks: semantic prediction and modality alignment. The semantic prediction task uses lightweight attention mechanisms ($O(|\mathcal{U}|(M_{\text{ex}}^{2}d_{t}+M_{\text{ex}}d_{t}))$) and scoring ($O(|\mathcal{U}|Nd_{t})$), while modality alignment employs contrastive learning ($O(|\mathcal{U}|(M_{\text{im}}d+M_{\text{ex}}d_{t}+Nd))$), both incurring minimal computational overhead while significantly improving generalization and recommendation accuracy.
Overall, the total time complexity of the training phase in EIMF is:
$O(I|\mathcal{U}|^{2}+E(|\mathcal{U}|Nd+|\mathcal{U}|Nd_{t}+P|\mathcal{U}|)),$
where the first term represents the one-time offline precomputation cost, and the second term grows linearly with the number of training epochs $E$. By combining offline precomputation with efficient behavioral modeling, EIMF achieves a favorable balance between computational cost and model effectiveness during training. During the serving phase, EIMF relies solely on the IBIM module for recommendation, without invoking the LLM module. The inference complexity remains at $O(|\mathcal{U}|Nd)$, ensuring low latency and high scalability.
In summary, EIMF fully leverages the semantic reasoning capabilities of LLMs during training while minimizing computational and time costs during inference, achieving an effective balance between performance enhancement and efficiency optimization. This design provides a practical and scalable solution for industrial-scale RSs.
@@ 5.1.1. Datasets and Evaluation Metrics
Datasets Users Items Interactions Avg.len. Sparsity(%)
Beauty 22,363 12,101 198,502 8.87 99.93
Grocery 14,681 8,713 151,254 10.30 99.88
Office 4,905 2,420 53,258 10.85 99.55
Ml-1M 6,040 3,416 99,9611 165.49 95.16
To evaluate the framework’s performance, we selected three sub-datasets from the review dataset of the Amazon e-commerce platform [^1] https://jmcauley.ucsd.edu/data/amazon/links.html, namely Office Products, Grocery & Gourmet Food, and All Beauty. To further validate the applicability and effectiveness of our method in non-e-commerce scenarios, we incorporate the MovieLens(Ml)-1M  [^2] https://grouplens.org/datasets/movielens/ dataset, which contains over 1 million ratings from more than 6,000 users on over 4,000 movies, featuring long user behavior sequences with rich continuity and diversity characteristics. The statistics of the four datasets after preprocessing are summarized in Table 3.
5.1.2. Evaluation Metrics
In terms of metrics, we chose the Recall @$K$, Normalized Discounted Cumulative Gain (NDCG) @$K$, and Hit Ratio (HR) @$K$, which are commonly used in SR. The value of $K$ is set to 20 and 50.
The calculation formula of Recall @$K$ is as follows:
$\text{Recall}@K=\frac{|\mathcal{R}_{K}\cap\mathcal{T}|}{|\mathcal{T}|}$
Here, $\mathcal{R}_{K}$ denotes the set of Top-$K$ items recommended by the system, $\mathcal{T}$ represents the set of ground-truth items interacted by the user in the test set, ||$\cdot$|| ⋅ | indicates the cardinality of a set (i.e., the number of elements).
The calculation formula of NDCG @$K$ is as follows:
$\text{DCG}@K=\sum_{i=1}^{K}\frac{\text{rel}_{i}}{\log_{2}(i+1)}$
$\text{IDCG}@K=\sum_{i=1}^{\min(|\mathcal{T}|,K)}\frac{\text{rel}_{i}^{\text{ideal}}}{\log_{2}(i+1)}$
$\text{NDCG}@K=\frac{\text{DCG@K}}{\text{IDCG@K}}$
In these equations, $\text{rel}_{i}$ represents the relevance score of the item at rank $i$ in the recommended list, typically binary in this paper (i.e., $\text{rel}_{i}=1$ for relevant items and $\text{rel}_{i}=0$ for irrelevant items), $\text{rel}_{i}^{\text{ideal}}$ is the relevance score of the item at rank $i$ in the ideal ranking sorted by relevance in descending order, $i$ denotes the rank position in the recommended list ranging from 1 to $K$, $\mathcal{T}$ is the set of ground-truth items interacted by the user in the test set, $|\mathcal{T}|$ indicates the number of relevant items in the test set, $\min(|\mathcal{T}|,K)$ ensures IDCG accounts for all relevant items up to $K$, $\log_{2}(i+1)$ is the discount factor reducing the contribution of items at lower ranks.
The calculation formula of HR @$K$ is as follows:
$\text{HR}@K=\frac{|\{u\in\mathcal{U}:\mathcal{R}_{K}^{u}\cap\mathcal{T}^{u}\neq\emptyset\}|}{|\mathcal{U}|}$
In this formula, $\mathcal{U}$ represents the set of all users, $u$ denotes a single user where $u\in\mathcal{U}$, $\mathcal{R}_{K}^{u}$ is the set of Top-$K$ items recommended for user $u$, $\mathcal{T}^{u}$ is the set of ground-truth items interacted by user $u$ in the test set, $\mathcal{R}_{K}^{u}\cap\mathcal{T}^{u}\neq\emptyset$ indicates that at least one item in the recommended list for user $u$ is relevant, ||$\cdot$|| ⋅ | denotes the cardinality of a set (i.e., the number of elements).
5.1.3. Implementation Details and Hyperparameter Settings
The Qwen-Turbo model (Bai et al., 2023) was chosen for the LLM in the experiments and is responsible for inference about user interests. All EIMFs not specifically marked use REMI as the backbone of the multi-interest SR model. All models were implemented in Pytorch 1.10.0 and Python 3.9 in the conda environment. We follow previous research(Xie et al., 2023), and in our experiments, the batch size $=128$, the dimension $=64$, and the maximum number of training iterations for all models is 1 million. The number of interests for the multi-interest model is set to 4, the number of user interests inferred by LLM to at most $20$ and the optimizer is trained using Adam with the learning rate set to $0.001$. Please note that, except for the hyperparameter experiments, we use $p$ = -10 as the reference value setting for the AP algorithm in the rest of the experiments (clustered typical samples account for about 3% of the dataset).
Table 4. The result of performance comparison between baselines and EIMF on three datasets.
Dataset Model Pop DNN GRU4Rec SASRec BERT4Rec LLM2BERT4Rec MIND ComiRec-SA REMI EIMF Improv.(%)
Grocery $\text{Recall}@20$ 0.0729 0.1252 0.1387 0.1536 0.1544 0.1259 0.1445 0.1122 0.1617 0.1733 +7.17
$\text{Recall}@50$ 0.1305 0.2044 0.2300 0.2508 0.2372 0.2080 0.2162 0.2076 0.2574 0.2964 +15.15
$\text{NDCG}@20$ 0.0448 0.0804 0.0905 0.1041 0.1074 0.0793 0.0844 0.0706 0.0953 0.1006 -6.33
$\text{NDCG}@50$ 0.0624 0.0969 0.1097 0.1139 0.1130 0.0954 0.0967 0.0918 0.1108 0.1233 +8.25
$\text{HR}@20$ 0.1252 0.2076 0.2321 0.2586 0.2614 0.2164 0.2328 0.1851 0.2539 0.2710 +4.80
$\text{HR}@50$ 0.2137 0.3227 0.3649 0.3750 0.3608 0.3294 0.3322 0.3220 0.3832 0.4234 +10.49
@@ 5.1.3. Implementation Details and Hyperparameter Settings
Office $\text{Recall}@20$ 0.0512 0.0866 0.0800 0.1152 0.1037 0.1085 0.0915 0.0788 0.1072 0.1222 +6.08
$\text{Recall}@50$ 0.0932 0.1764 0.1770 0.1963 0.1920 0.2117 0.1593 0.1589 0.2018 0.2296 +8.45
$\text{NDCG}@20$ 0.0331 0.0509 0.0537 0.0643 0.0635 0.0660 0.0513 0.0456 0.0637 0.0649 -1.66
$\text{NDCG}@50$ 0.0472 0.0697 0.0809 0.0805 0.0834 0.0939 0.0706 0.0653 0.0825 0.0930 -0.95
$\text{HR}@20$ 0.0713 0.1466 0.1384 0.1792 0.1812 0.1812 0.1466 0.1202 0.1751 0.1772 -2.20
$\text{HR}@50$ 0.1426 0.2770 0.2811 0.2871 0.3116 0.3360 0.2505 0.2321 0.2973 0.3503 +4.25
@@
  • 2
    *The best baseline value is underlined, and the best performance among all models is in bold.
(a) Recall@$K$
(b) NDCG@$K$
Figure 3. The results of different $K$ on the Ml-1M dataset
@@ 5.1.4. Baselines
Regarding baseline selection, we mainly select from single-interest and multi-interest aspects, including single-interest models (Covington et al., 2016; Hidasi, 2015; Kang and McAuley, 2018; Sun et al., 2019), LLM-based RS model (Boz et al., 2024), and multi-interest models (Li et al., 2019; Cen et al., 2020; Xie et al., 2023).
  • Pop. An algorithm that makes recommendations based on item popularity.
  • DNN (Covington et al., 2016). The twin-tower DNN model developed by the YouTube team pools user behaviors and then uses an MLP to obtain user interest representations.
@@
  • BERT4Rec (Sun et al., 2019). An SR model based on the BERT architecture, which uses a PLM to process user behavior sequences and captures contextual information in the sequence through a bidirectional encoder.
@@
  • LLM2BERT4Rec (Boz et al., 2024). A framework to enhance existing SR models by leveraging semantically rich item representations provided by LLM.
@@ 5.2.1. Performance Comparison Experiment with Baselines
We added the EIMF framework to REMI to conduct performance comparison experiments with other baselines. The results are shown in Table 4. Here we marked the best performance value among all models in bold and the best performance value in the baseline with an underline. Based on the results in Table 4, we can make the following observations:
@@
  • In the Grocery and Beauty datasets in Table 4, the performance of the multi-interest model is significantly better than that of the single-interest model, and as the amount of data increases, this advantage becomes more and more significant. Our analysis shows that multi-interest learning is more suitable for data-rich environments because it is able to capture more complex and diverse underlying interests among users or items, thereby more accurately understanding subtle differences between individuals.
@@
  • In the Office dataset shown in Table 4, the traditional single-interest SR model outperforms the multi-interest model. Our analysis shows that when the amount of data is limited, it is difficult to accurately capture the diverse interests of users by simply relying on user behavior information. Therefore, a model that focuses on a single major interest can utilize the limited information in a more concentrated manner, thus providing more accurate recommendation results.
@@
  • In the large datasets (i.e., Grocery and Beauty), EIMF consistently achieved the best performance with a large margin compared to the baselines, while in the small dataset (Office), EIMF achieved the second-best performance, which slightly falls behind LLM2BERT4Rec. First, this shows that the rich semantic knowledge in the LLM can effectively enhance the behavioral modeling ability of user representation, showing the great potential of LLM in recommendation tasks. Second, although some metrics of EIMF are slightly lower than LLM2BERT4Rec on small datasets, its better performance on large datasets proves that EIMF not only has good stability but also can adapt to more diverse data environments.
5.2.2. Performance Experiment on Different Top-$K$ Metrics
To comprehensively evaluate the effectiveness of the EIMF framework, we conducted experiments to assess its performance across various Top-$K$ metrics. We use REMI as the base model and test the performance changes of K=5, 10, and 20 in the Ml-1M dataset. Figure 3 illustrates that EIMF consistently outperforms the original REMI model across all Top-$K$ metrics on the Ml-1M dataset. Specifically, Recall@20 improves from 0.0834 to 0.0898 (a 7.7% increase), and NDCG@20 rises from 0.2851 to 0.2996 (a 5.1% increase). For smaller K values, the gains are slightly more pronounced: Recall@5 increases from 0.0175 to 0.0203 (16.0% improvement), and NDCG@5 improves from 0.1410 to 0.1749 (24.0% improvement).
We attribute the performance gains achieved by incorporating EIMF to its core module, ESIM, which leverages LLMs to infer semantic-level user interests from textualized interaction histories. The semantic augmentation offered by EIMF is particularly effective for Top-$5$ recommendation metrics, as it enables fine-grained interest modeling to address the data sparsity issue. For instance, a user with sparse ratings for environmental documentaries may still be semantically linked, via ESIM, to a cohort of users who are passionate about this genre. This allows EIMF to infer the user’s preference for “ocean conservation documentaries” and recommend highly relevant items, such as The Cove. In contrast, Top-$20$ metrics typically require less precision, as broader genre-based recommendations (e.g., trending documentaries) may suffice. Such nuanced preferences are often underrepresented in purely implicit feedback signals, and thus, the semantic modeling enhances the granularity of interest representation. Additionally, EIMF employs a prototypical sampling strategy that captures group-level interest patterns while significantly reducing the inference overhead of the LLM. This approach is especially well-suited to dense datasets like ML-1M, where user behaviors exhibit strong intra-cluster similarity, enabling ESIM to generalize semantic insights across users effectively.
Table 5. Performance experimental results of different backbones on Beauty, Office and Ml-1M datasets.
Dataset Beauty Office Ml-1M
Model $\text{Recall}@20$ $\text{NDCG}@20$ $\text{HR}@20$ $\text{Recall}@20$ $\text{NDCG}@20$ $\text{HR}@20$ $\text{Recall}@20$ $\text{NDCG}@20$ $\text{HR}@20$
BERT4Rec 0.1405 0.0780 0.2096 0.1037 0.0635 0.1812 0.1294 0.3643 0.7698
EIMF(BERT4Rec) 0.1421 0.0804 0.2181 0.1186 0.0719 0.1874 0.1405 0.4011 0.8228
Improv.(%) +1.14 +3.08 +4.06 +14.36 +17.11 +3.42 +8.58 +10.10 +6.88
MIND 0.1563 0.0787 0.2204 0.0915 0.0513 0.1466 0.0605 0.2450 0.5977
EIMF(MIND) 0.1755 0.0896 0.2494 0.0930 0.0499 0.1446 0.0654 0.2496 0.6043
Improv.(%) +5.62 +7.55 +4.64 +26.91 +12.16 +17.88 +8.10 +1.88 +1.09
REMI 0.2189 0.1139 0.3111 0.1072 0.0636 0.1751 0.0834 0.2851 0.6804
EIMF(REMI) 0.2322 0.1203 0.3268 0.1222 0.0646 0.1772 0.0897 0.2996 0.6854
Improv.(%) +6.31 +3.37 +5.95 +8.45 -0.95 +4.25 +7.55 +5.09 +0.73